The following sample of eight measurements was randomly selected from a normally distributed population: 12, 10, 9, 8, 15, 10, 11, and 13. Test for significant difference between the sample mean and the population mean of 10. Use 𝛼 = 0.05.


1
Expert's answer
2021-11-08T21:20:55-0500

First determine the value of the sample mean and standard deviation.

The mean is given by,

"\bar{x}=\sum(x)/n" where "n=8"

"\bar{x}=88/8=11"

To find the standard deviation, we need to find the variance given by the formula,

"V(X)=(\sum(x^2)-(\sum(x))^2/n)/(n-1)"

"V(X)=(1004-968)\/7=5.142857"

The standard deviation "SD(X)=\sqrt{V(X)}=\sqrt{5.142857}=2.267787"

The hypotheses tested are,

"H_0:\mu=10\space Against\space H_1:\mu\not=10"

To perform this test, we shall apply the t distribution since the population variance is not known and the sample size is less than 30.

The test statistic is given as,

"t=(\bar{x}-\mu)/(SD(X)/\sqrt{n})"

"t=(11-10)/(2.267787/\sqrt{8})=1/0.801784=1.247219"

"t" is compared with the table value at "\alpha=0.05" with "(n-1)=8-1=7" degrees of freedom.

Thus, "t_{\alpha/2,7}=t_{0.05/2,7}=t_{0.025,7}=2.365", and the null hypothesis is rejected if "t\gt t_{0.025,7}"

Since "t=1.247219\lt t_{0.025,7}=2.365", we fail to reject the null hypothesis and we conclude that there is no sufficient evidence to show that the sample mean and the population mean are significantly different at 5% significance level.